# Exploring Homogeneous and Heterogeneous Consistent Label Associations for Unsupervised Visible-Infrared Person ReID

### v2

@@ 3.1 Modality-Unified Label Transfer
$\mathcal{I}^{r}_{he}(\hat{\mathbf{y}}^{r})=\sum_{i=1}^{N^{r}}\sum_{j=1}^{N^{v}}\mathbf{S}^{he(rv)}_{ij}\|\hat{\mathbf{y}}^{r}_{i}-\tilde{\mathbf{y}}_{i}^{v}\|^{2}_{2}.$
Nonetheless, solely minimizing the above two inconsistency terms can lead to a collapse, where all instances end up with nearly identical labels. This is because these two inconsistency statements do not consider the fact that pairwise instances with low affinities should have distinct labels. To avoid such collapse, we propose the self-inconsistency to constraint the Euclidean distance between the transferred labels and the initial coarse labels:
$\mathcal{I}^{e}_{self}({\mathbf{y}}^{e})=\sum_{i=1}^{N^{e}}\|\mathbf{y}_{i}^{e}-{\mathbf{y}}_{i}^{e}(0)\|^{2}_{2}.\vspace{-1mm}$ (9)
We minimize the above three terms of inconsistency, aiming to derive the pseudo-labels $\mathbf{y}^{e}$ with the lowest inconsistency. This optimization can be formulated as follows:
$\min\limits_{\mathbf{y}^{e}}\quad\mathcal{I}^{e}_{ho}(\mathbf{y}^{e})+\alpha\mathcal{I}^{e}_{self}(\mathbf{y}^{e})+(1-\alpha)\mathcal{I}^{e}_{he}(\mathbf{y}^{e}),\\ \vspace{-1mm}$ (10)
@@
Label Transfer. The above optimization problem can be solved by the LGC algorithm [58]. Specifically, the visible and infrared modality-unified labels are updated alternately as Eq.11 and Eq.12 until convergence:
@@
$\begin{array}[]{lc}\vspace{1ex}\tilde{\mathbf{y}}^{v}(t+1)=(1-\alpha)\mathbf{S}^{he(vr)}\cdot\hat{\mathbf{y}}^{r}(t)+\alpha\tilde{\mathbf{y}}^{v}(0),\\ \vspace{1ex}\tilde{\mathbf{y}}^{v}(t+1)=\frac{1}{2}\left(\mathbf{S}^{ho(v)}\cdot\tilde{\mathbf{y}}^{v}(t+1)+\tilde{\mathbf{y}}^{v}(t+1)\right),\end{array}\vspace{-6mm}$ (11)
@@
$\begin{array}[]{lc}\vspace{1ex}\hat{\mathbf{y}}^{r}(t+1)=(1-\alpha)\mathbf{S}^{he(rv)}\cdot\tilde{\mathbf{y}}^{v}(t)+\alpha\hat{\mathbf{y}}^{r}(0),\\ \vspace{1ex}\hat{\mathbf{y}}^{r}(t+1)=\frac{1}{2}\left(\mathbf{S}^{ho(r)}\cdot\hat{\mathbf{y}}^{r}(t+1)+\hat{\mathbf{y}}^{r}(t+1)\right),\end{array}\vspace{-2mm}$ (12)
where $\mathbf{y}^{e}(t)=\{\tilde{\mathbf{y}}^{v}(t),\hat{\mathbf{y}}^{r}(t)\}$ denotes the labels in $t-th$ iteration during the optimization process.
Nevertheless, the transferred pseudo-labels are excessively smoothed, which is detrimental to the network’s ability to learn predictions with low entropy. Therefore, we derive the soft modality-unified labels by aggregating both hard labels and soft labels, which can be formulated as:
$\mathbf{y}^{e}=\beta\cdot\mathbf{y}_{hard}^{e}+(1-\beta)\cdot\mathbf{y}_{soft}^{e},\vspace{0mm}$ (13)
where $\mathbf{y}_{hard}$ and $\mathbf{y}_{soft}$ denotes the hard pseudo-labels and the soft pseudo-labels obtained from the transferred labels $\mathbf{y}^{e}=\{\tilde{\mathbf{y}}^{v},\hat{\mathbf{y}}^{r}\}$, respectively. $\beta$ is a parameter to control the smoothness of the pseudo-label. It is noteworthy that the labels $\tilde{\mathbf{y}}_{i}^{v}$ and $\hat{\mathbf{y}}_{i}^{r}$ have a same dimension of $K$.
3.2 Modality-Invariant Representation Learning
Guided by the modality-unified soft labels from the MULT module, we lay emphasis on alleviating the cross-modality discrepancy while keeping intra-modality consistency. Following the MSMA framework [5], we concurrently execute intra-modality and cross-modality contrastive learning. To be specific, we construct a cross-modality memory bank $\mathbf{M}^{a}\in\mathbb{R}^{K^{v}\times d}$ and an intra-cross-modality memory bank $\hat{\mathbf{M}}^{r}\in\mathbb{R}^{K^{v}\times d}$, which are initialized in the same manner as the visible intra-modality memory bank $\tilde{\mathbf{M}}^{v}$. Following the widely-used memory-based methods [15, 11, 47, 7], we carry out contrastive learning during forward-propagation (FP) and update the memory banks during backward-propagation (BP). The visible intra-modality contrastive loss can be formulated as:
@@ 3.2 Modality-Invariant Representation Learning
$\mathcal{L}^{v}_{IM}=\frac{1}{B}\sum^{B}_{i=1}\mathcal{L}_{ce}(\mathbf{P}(\mathbf{f}_{i}^{v}|\tilde{\mathbf{M}}^{v},\tau),\tilde{\mathbf{y}}^{v}_{i}),\vspace{-1mm}$ (14)
@@
where $\mathcal{L}_{ce}$ denotes the standard cross-entropy loss. For infrared modality, every instance owns two types of labels, i.e., $\tilde{\mathbf{y}}^{r}$ and $\hat{\mathbf{y}}^{r}$. Such two types of labels are utilized jointly for infrared intra-modality contrastive learning, which can be formulated as:
@@
These two types of labels play a mutual corrective role and facilitate intra-modality training. The total intra-modality contrastive loss can be formulated as follows:
$\mathcal{L}_{IM}=\mathcal{L}^{v}_{IM}+\mathcal{L}^{r}_{IM}.\vspace{-0.5mm}$ (16)
Besides, we perform cross-modality contrastive learning to learn modality-invariant representations:
$\mathcal{L}_{CM}$ $=\frac{1}{B}\sum^{B}_{i=1}\mathcal{L}_{ce}(\mathbf{P}(\mathbf{f}_{i}^{v}|\mathbf{M}^{a},\tau),\tilde{\mathbf{y}}^{v}_{i})$ (17)
$+\frac{1}{B}\sum^{B}_{i=1}\mathcal{L}_{ce}(P(\mathbf{f}_{i}^{r}|\mathbf{M}^{a},\tau),\hat{\mathbf{y}}^{r}_{i}).$
@@
During the BP stage, the memory banks are updated with the input features in a momentum strategy [18, 11]:
$\mathbf{m}(y_{i}^{e})\leftarrow\mu\mathbf{m}(y_{i}^{e})+(1-\mu)\mathbf{f}(y_{i}^{e}),\vspace{-0.5mm}$ (18)
@@
where $\mu$ is the momentum updating factor. $\mathbf{m}(y_{i}^{e})$ is the $y_{i}^{e}-th$ prototype in the memory bank $\mathbf{M}$, and $\mathbf{f}(y_{i}^{e})$ is the input feature with label $y_{i}^{e}$ in current mini-batch. Note that the hard label $y_{i}^{e}$ for memory update is derived from the soft modality-unified label $\mathbf{y}_{i}^{e}$.
@@
Table 1: Comparison with the state-of-the-art methods on SYSU-MM01 and RegDB. ‘${}^{\ast}$’ denotes GUR without camera labels. Since our method does not require any camera label information, for fair comparison we do not report the results of GUR with camera labels.

### v4

@@ 3.1 Modality-Unified Label Transfer
Nonetheless, solely minimizing the above two inconsistency terms can lead to a collapse, where all instances end up with nearly identical labels. This is because these two inconsistency statements do not consider the fact that pairwise instances with low affinities should have distinct labels. Such a reduction in label diversity can negatively impact network training. To avoid such collapse, a label initialization approach that ensures the diversity of the pseudo-labels is necessary.
Specifically, the intra-modality labels are initialized according to the cluster centroids, and the cross-modality labels are initialized by the widely-used OTLA Wang et al (2022) method. To fully utilize the relationships between instances and each cluster centroid, we use the soft probability distribution from the corresponding memory bank as initial visible intra-modality labels $\tilde{\mathbf{y}}^{v}$:
$\tilde{\mathbf{y}}^{v}\in\mathbb{R}^{N^{v}\times K^{v}},\tilde{\mathbf{y}}^{v}_{i}(0)=\mathbf{P}(\mathbf{f}_{i}^{v}|\tilde{\mathbf{M}}^{v},\tau)\in\mathbb{R}^{K^{v}},$ (7)
${{P}_{j}(\mathbf{f}|\mathbf{M},\tau)}=\frac{\exp\left(\mathbf{f}^{\top}\mathbf{c}_{j}/\tau\right)}{\sum_{k=1}^{K}\exp\left(\mathbf{f}^{\top}\mathbf{c}_{k}/\tau\right)},$ (8)
$\mathbf{P}(\mathbf{f}|\mathbf{M},\tau)=[P_{1}(\mathbf{f}|\mathbf{M},\tau),\cdots,P_{K}(\mathbf{f}|\mathbf{M},\tau)],$ (9)
where $K$ denotes the number of prototypes in the memory $\mathbf{M}$, $\mathbf{c}_{j}$ and $\mathbf{c}_{k}$ represents $j$-th and $k$-th cluster prototypes in $\mathbf{M}$. Following Wang et al (2022); Cheng et al (2023b), we utilize Optimal Transport Label Assignment (OTLA) to initialize infrared cross-modality labels $\hat{\mathbf{y}}^{r}$:
$\mathop{\min}\limits_{\mathbf{P}}\langle\mathbf{P},\mathbf{C}\rangle+\frac{1}{\lambda}\langle\mathbf{P},-\log(\mathbf{P})\rangle.$ (10)
$s.t.\quad\mathbf{P}\mathds{1}=\mathds{1}\cdot\frac{1}{N^{r}},\quad\mathbf{P}^{\top}\mathds{1}=\mathds{1}\cdot\frac{1}{K^{v}},$
where $N^{r}$ denotes the number of infrared instances, and $K^{v}$ denotes the number of visible clusters. $\mathbf{C}\in\mathbb{R}^{N^{r}\times K^{v}}$ represents the cost matrix and $\mathbf{P}\in\mathbb{R}^{N^{r}\times K^{v}}$ represents the transport plan. We utilize the Euclidean distance between infrared instances and visible prototypes from $\tilde{\mathbf{M}}^{v}$ to calculate the cost, where $\mathbf{C}_{ij}=\|\mathbf{f}^{r}_{i}-\tilde{\mathbf{m}}^{v}_{j}\|_{2}^{2}$. $\mathbf{f}^{r}_{i}$ is the $i$-th feature vector in training set and $\tilde{\mathbf{m}}^{v}_{j}$ is the j-th prototype in memory $\tilde{\mathbf{M}}^{v}$. The initialized $\hat{\mathbf{y}}^{r}(0)\in\mathbb{R}^{N^{r}\times K^{v}}$ is the one-hot encoding form of the optimal solution $\mathbf{P}^{\ast}$ of Eq.10, which is formulated as follows:
${{\hat{\mathbf{y}}^{r}_{ik}(0)=\left\{\begin{array}[]{ll}1,\quad\text{if}\quad k=\arg\max_{j}\mathbf{P}^{\ast}_{ij}\\ 0,\quad\text{otherwise}\end{array},\right.\quad\hat{\mathbf{y}}^{r}_{i}(0)\in\mathbb{R}^{K^{v}}.}}$ (11)
In the optimization process, we aim to ensure that the transferred labels do not deviate too far from the initial labels, thereby guaranteeing diversity in the transferred labels. Therefore, we propose the self-inconsistency to constraint the Euclidean distance between the transferred labels and the initial coarse labels:
$\mathcal{I}^{v}_{self}({\tilde{\mathbf{y}}}^{v})=\sum_{i=1}^{N^{v}}\|\tilde{\mathbf{y}}_{i}^{v}-{\tilde{\mathbf{y}}}_{i}^{v}(0)\|^{2}_{2}.$ (12)
$\mathcal{I}^{r}_{self}({\hat{\mathbf{y}}}^{r})=\sum_{i=1}^{N^{r}}\|\hat{\mathbf{y}}_{i}^{r}-{\hat{\mathbf{y}}}_{i}^{r}(0)\|^{2}_{2}.$ (13)
We minimize the above three terms of inconsistency, aiming to derive the pseudo-labels $\tilde{\mathbf{y}}^{v}$ and $\hat{\mathbf{y}}^{r}$ with the lowest inconsistency. This optimization can be formulated as follows:
$\min\limits_{\tilde{\mathbf{y}}^{v}}\quad\mathcal{I}^{e}_{ho}(\tilde{\mathbf{y}}^{v})+\alpha\mathcal{I}^{e}_{self}(\tilde{\mathbf{y}}^{v})+(1-\alpha)\mathcal{I}^{e}_{he}(\tilde{\mathbf{y}}^{v}),\$ (14)
$\min\limits_{\hat{\mathbf{y}}^{r}}\quad\mathcal{I}^{e}_{ho}(\hat{\mathbf{y}}^{r})+\alpha\mathcal{I}^{e}_{self}(\hat{\mathbf{y}}^{r})+(1-\alpha)\mathcal{I}^{e}_{he}(\hat{\mathbf{y}}^{r}),\$ (15)
@@
Label Transfer. The above optimization problem can be solved by the LGC algorithm Zhou et al (2003). Specifically, the visible and infrared modality-unified labels are updated alternately as Eq.16 and Eq.17 until convergence:
@@
$\begin{array}[]{lc}\vspace{1ex}\tilde{\mathbf{y}}^{v}(t+1)^{\ast}=(1-\alpha)\mathbf{S}^{he(vr)}\cdot\hat{\mathbf{y}}^{r}(t)+\alpha\tilde{\mathbf{y}}^{v}(0),\\ \tilde{\mathbf{y}}^{v}(t+1)=\frac{1}{2}\mathbf{S}^{ho(v)}\cdot\tilde{\mathbf{y}}^{v}(t+1)^{\ast}+\frac{1}{2}\tilde{\mathbf{y}}^{v}(t+1)^{\ast},\end{array}$ (16)
@@
$\begin{array}[]{lc}\vspace{1ex}\hat{\mathbf{y}}^{r}(t+1)^{\ast}=(1-\alpha)\mathbf{S}^{he(rv)}\cdot\tilde{\mathbf{y}}^{v}(t)+\alpha\hat{\mathbf{y}}^{r}(0),\\ \vspace{1ex}\hat{\mathbf{y}}^{r}(t+1)=\frac{1}{2}\mathbf{S}^{ho(r)}\cdot\hat{\mathbf{y}}^{r}(t+1)^{\ast}+\frac{1}{2}\hat{\mathbf{y}}^{r}(t+1)^{\ast},\end{array}\vspace{-2mm}$ (17)
where $\mathbf{y}^{e}(t)=\{\tilde{\mathbf{y}}^{v}(t),\hat{\mathbf{y}}^{r}(t)\}$ denotes the labels in $t-th$ iteration during the optimization process. $\mathbf{y}^{e}(t+1)^{\ast}$ denotes the intermediate variables during the update process in one iteration. The upper terms in Eq.16 and Eq.17 can be regarded as a cross-modality transfer process, while the lower terms stand for a manner to further compensate the transferred pseudo-labels utilizing intra-modality structural information.
In detail, take the pseudo-labels $\tilde{\mathbf{y}}^{v}$ in visible modality, the inconsistency formulation Eq.14 can be expressed in the form of vectors and matrices:
$\mathcal{I}(\tilde{\mathbf{y}}^{v})$ $=\alpha\mathbf{tr}((\tilde{\mathbf{y}}^{v}-\tilde{\mathbf{y}}^{v}(0))^{\top}(\tilde{\mathbf{y}}^{v}-\tilde{\mathbf{y}}^{v}(0)))$ (18)
$+(1-\alpha)\mathbf{tr}((\tilde{\mathbf{y}}^{v}-\mathbf{\mathbf{S}}^{he(vr)}\cdot\hat{\mathbf{y}}^{r})^{\top}(\tilde{\mathbf{y}}^{v}-\mathbf{\mathbf{S}}^{he(vr)}\cdot\hat{\mathbf{y}}^{r}))$
$+\mathbf{tr}(\tilde{\mathbf{y}}^{v\top}(\mathbf{I}_{N^{v}}-\mathbf{S}^{ho(v)}){\tilde{\mathbf{y}}^{v}}),$
where $\mathbf{I}_{N^{v}}$ denotes the $N^{v}\times N^{v}$ identity matrix and $\mathbf{tr}(\cdot)$ denotes the trace of matrix. To minimize the inconsistency $\mathcal{I}(\tilde{\mathbf{y}}^{v})$ for visible pseudo-labels, we compute its partial derivative with respect to $\tilde{\mathbf{y}}^{v}$ and equate this partial derivative to zero, where $\frac{\partial{\mathcal{I}(\tilde{\mathbf{y}}^{v})}}{\partial{\tilde{\mathbf{y}}^{v}}}=0$ Then $\tilde{\mathbf{y}}^{v}$ can be represented as:
@@
${\tilde{\mathbf{y}}^{v}=\frac{1}{2}\mathbf{S}^{ho(v)}\cdot\tilde{\mathbf{y}}^{v}+\frac{1}{2}\left[(1-\alpha)\mathbf{S}^{he(vr)}\cdot\hat{\mathbf{y}}^{r}+\alpha\tilde{\mathbf{y}}^{v}(0)\right].}\vspace{0mm}$ (19)
To enhance its convergence stability, we decompose Eq.19 into two steps, $i.e.$, cross-modality transfer, and intra-modality transfer, which is formulated as follows:
[…]
During each iterative optimization process, the pseudo-labels $\tilde{\mathbf{y}}^{v}$ are updated by alternating iterations of fixed points, which is formulated as Eq.16.
Input: $\mathbf{f}^{v}$: Visible features; $\mathbf{f}^{r}$: Infrared features;
Output:
Visible intra-modality labels $\tilde{\mathbf{y}}^{v}$;
Infrared cross-modality labels $\hat{\mathbf{y}}^{r}$;
Initialization:
Two intra-modality memory banks $\tilde{\mathbf{M}}^{v}$ and $\tilde{\mathbf{M}}^{r}$;
Intra-modality visible labels $\tilde{\mathbf{y}}^{v}(0)$ (Eq.7);
Cross-modality infrared labels $\hat{\mathbf{y}}^{r}(0)$ (Eq.10);
Affinity Modeling:
Intra-modality affinity matrices $\mathbf{S}^{v}$ and $\mathbf{S}^{r}$
(Eq.1 & Eq.2);
Cross-modality affinity matrices $\mathbf{S}^{vr}$ and $\mathbf{S}_{rv}$ (Eq.3);
Do Label Transfer (V2R):
$t=0$; $\epsilon_{0}=1e-2$; $\epsilon=1e6$;
while $\epsilon>\epsilon_{0}$ do
      Calculate $\tilde{\mathbf{y}}^{v}(t+1)$ according to Eq.16;
       Calculate $\hat{\mathbf{y}}^{r}(t+1)$ according to Eq.17;
       ϵitalic-$\epsilon\leftarrow$
       $\max(\|\tilde{\mathbf{y}}^{v}(t+1)-\tilde{\mathbf{y}}^{v}(t)\|_{1},\|\hat{\mathbf{y}}^{r}(t+1)-\hat{\mathbf{y}}^{r}(t)\|_{1})$;
       $t\leftarrow t+1$;
Update labels $\tilde{\mathbf{y}}^{v}\leftarrow\tilde{\mathbf{y}}^{v}(t)$; $\hat{\mathbf{y}}^{r}\leftarrow\hat{\mathbf{y}}^{r}(t)$;
Return: Soft labels $\tilde{\mathbf{y}}^{v}$ and $\hat{\mathbf{y}}^{r}$ (Eq.21).
Algorithm 1 V2R MULT
Nevertheless, the transferred pseudo-labels are excessively smoothed, which is detrimental to the network’s ability to learn predictions with low entropy. Therefore, we derive the soft modality-unified labels by aggregating both hard labels and soft forms of the pseudo-labels, which can be formulated as:
$\mathbf{y}^{e}=\beta\cdot\mathbf{y}_{hard}^{e}+(1-\beta)\cdot\mathbf{y}_{soft}^{e},$ (21)
where $\beta$ is a parameter to control the smoothness of the pseudo-label. $\mathbf{y}^{e}_{hard}$ and $\mathbf{y}^{e}_{soft}$ denotes the hard one-hot encoding pseudo-labels and the soft pseudo-labels obtained from the transferred labels $\mathbf{y}^{e}=\{\tilde{\mathbf{y}}^{v},\hat{\mathbf{y}}^{r}\}$, respectively. The one-hot form $\mathbf{y}^{e}_{hard}$ is directly derived from the soft label $\mathbf{y}^{e}_{soft}$.
It is noteworthy that, the labels $\tilde{\mathbf{y}}_{i}^{v}$ and $\hat{\mathbf{y}}_{i}^{r}$ have a same dimension of $K^{v}$. The detailed algorithm flow of our V2R MULT is in Alg.1. The R2V MULT, which provides $\tilde{\mathbf{y}}^{r}$ and $\hat{\mathbf{y}}^{v}$ in infrared label space, holds a symmetrical form with V2R MULT.
3.2 Alternative Modality-Invariant Representation Learning
Guided by the modality-unified soft labels from the MULT module, we lay emphasis on alleviating the cross-modality discrepancy while keeping intra-modality consistency. We concurrently execute intra-modality and cross-modality contrastive learning. Additionally, we design an alternative scheme that alternates between V-based and R-based training modes.
To be specific, for V-based training mode, we construct a cross-modality memory bank $\mathbf{M}^{c}\in\mathbb{R}^{K^{v}\times d}$ and an auxiliary memory bank $\hat{\mathbf{M}}^{r}\in\mathbb{R}^{K^{v}\times d}$, which are initialized in the same manner as the visible intra-modality memory bank $\tilde{\mathbf{M}}^{v}$. $\mathbf{M}^{c}$ is updated by instances from both modalities and $\hat{\mathbf{M}}^{r}$ is updated by instances only from the infrared modality. The auxiliary memory bank $\hat{\mathbf{M}}^{r}$ aims to learn the infrared intra-modality structure depicted by its labels $\hat{\mathbf{y}}^{r}$ in visible label space, thereby serving as a compensatory mechanism for $\tilde{\mathbf{M}}^{r}$. In fact, instances from different identities are sometimes gathered in the same cluster with their intra-modality labels ($\tilde{\mathbf{y}}^{r}$), while they are distributed correctly into different clusters with their cross-modality labels ($\hat{\mathbf{y}}^{r}$). Thus the auxiliary memory, which reflects the structure of $\hat{\mathbf{y}}^{r}$, can assist the intra-modality learning. As shown in Flg.3, compared to utilizing prototypes solely from the intra-modality memory, leveraging prototypes from the auxiliary memory of another modality is advantageous for uncovering fine-grained structural details within clusters in both two modalities.
Figure 3: Illustration of the role of the auxiliary memory. Different shapes denote different identities and different edge colors denote different cross-modality labels.
Following the widely-used memory-based methods Ge et al (2020b); Dai et al (2023); Yang et al (2022); Cheng et al (2023c), we carry out contrastive learning during forward-propagation (FP) and update the memory banks during backward-propagation (BP). The visible intra-modality contrastive loss can be formulated as:
$\mathcal{L}^{v}_{IM}=\frac{1}{B}\sum^{B}_{i=1}\mathcal{L}_{ce}(\mathbf{P}(\mathbf{f}_{i}^{v}|\tilde{\mathbf{M}}^{v},\tau),\tilde{\mathbf{y}}^{v}_{i}),$ (22)
@@ 3.2 Alternative Modality-Invariant Representation Learning
where $\mathcal{L}_{ce}$ denotes the standard cross-entropy loss and $B$ denotes the input batch size. For infrared modality in V-based training, both $\tilde{\mathbf{y}}^{r}$ and $\hat{\mathbf{y}}^{r}$ are utilized jointly during contrastive learning with their corresponding memory $\tilde{\mathbf{M}}^{r}$ and $\hat{\mathbf{M}}^{r}$:
@@
These two types of labels in two different label spaces play a mutual corrective role and facilitate intra-modality training. The total intra-modality contrastive loss can be formulated as follows:
$\mathcal{L}_{IM}=\mathcal{L}^{v}_{IM}+\mathcal{L}^{r}_{IM}.$ (24)
Besides, we perform cross-modality contrastive learning to constrain the relationships between instances in different modalities and their corresponding modality-agnostic cluster prototypes in $\mathbf{M}^{c}$ in the feature space:
$\mathcal{L}_{CM}$ $=\frac{1}{B}\sum^{B}_{i=1}\mathcal{L}_{ce}(\mathbf{P}(\mathbf{f}_{i}^{v}|\mathbf{M}^{c},\tau),\tilde{\mathbf{y}}^{v}_{i})$ (25)
$+\frac{1}{B}\sum^{B}_{i=1}\mathcal{L}_{ce}(P(\mathbf{f}_{i}^{r}|\mathbf{M}^{c},\tau),\hat{\mathbf{y}}^{r}_{i}).$
@@
During the BP stage, the memory banks are updated with the input features in a momentum strategy He et al (2020); Dai et al (2023):
$\tilde{\mathbf{m}}^{v}(\tilde{y}_{i}^{v})\leftarrow\mu\tilde{\mathbf{m}}^{v}(\tilde{y}_{i}^{v})+(1-\mu)\mathbf{f}(\tilde{y}_{i}^{v}),\mathbf{f}\in\{\mathbf{f}_{1}^{v},\cdots,\mathbf{f}_{B}^{v}\}$ (26)
$\tilde{\mathbf{m}}^{r}(\tilde{y}_{i}^{r})\leftarrow\mu\tilde{\mathbf{m}}^{v}(\tilde{y}_{i}^{r})+(1-\mu)\mathbf{f}(\tilde{y}_{i}^{r}),\mathbf{f}\in\{\mathbf{f}_{1}^{r},\cdots,\mathbf{f}_{B}^{r}\}$
$\mathbf{m}^{c}(\tilde{y}_{i}^{v})\leftarrow\mu\mathbf{m}^{c}(\tilde{y}_{i}^{v})+(1-\mu)\mathbf{f}(\tilde{y}_{i}^{v}),\mathbf{f}\in\{\mathbf{f}_{1}^{v},\cdots,\mathbf{f}_{B}^{v}\}$
$\mathbf{m}^{c}(\hat{y}_{i}^{r})\leftarrow\mu\mathbf{m}^{c}(\hat{y}_{i}^{r})+(1-\mu)\mathbf{f}(\hat{y}_{i}^{r}),\mathbf{f}\in\{\mathbf{f}_{1}^{r},\cdots,\mathbf{f}_{B}^{r}\}$
$\hat{\mathbf{m}}^{r}(\hat{y}_{i}^{r})\leftarrow\mu\hat{\mathbf{m}}^{r}(\hat{y}_{i}^{r})+(1-\mu)\mathbf{f}(\hat{y}_{i}^{r}),\mathbf{f}\in\{\mathbf{f}_{1}^{r},\cdots,\mathbf{f}_{B}^{r}\},$
@@
where $\mu$ is the momentum updating factor. $\mathbf{m}(y)$ is the $y-th$ prototype in the memory bank $\mathbf{M}$, and $\mathbf{f}(y)$ is the input feature with label $y$ in current mini-batch. Note that $y=\{\tilde{y}_{i}^{v},\tilde{y}_{i}^{r},\hat{y}_{i}^{r}\}$ for memory update is the hard label form of the soft labels $\mathbf{y}=\{\tilde{\mathbf{y}}_{i}^{v},\tilde{\mathbf{y}}_{i}^{r},\hat{\mathbf{y}}_{i}^{r}\}$.
For R-based training, a cross-modality memory bank $\mathbf{M}^{c}\in\mathbb{R}^{K^{r}\times d}$ is constructed according to labels $\tilde{\mathbf{y}}^{r}$ and $\hat{\mathbf{y}}^{v}$ from R2V MULT. Meanwhile an auxiliary memory bank $\hat{\mathbf{M^{v}}}\in\mathbb{R}^{K^{r}\times d}$ is initialized in the same manner as $\tilde{\mathbf{M}}^{r}$. It is updated by visible features based on $\hat{\mathbf{y}}^{v}$ in the infrared label space. The loss functions for R-based training can be written in a symmetric form in comparison to the V-based training.
The entire training process is described in Alg.2. We conduct an alternate scheme for the V-based training and the R-based training as the epoch goes on: when epoch-id $E\%2==0$, the V-based training is conducted; otherwise, the R-based training is conducted. In one epoch, the model is supervised by labels from MULT of a specific direction. In adjacent epochs, the model is supervised by labels from MULT of different directions. Such an alternative scheme mitigates the inconsistency between pseudo-labels from MULT of different directions (V2R / R2V).
Input: epoch $E$; network $f_{\theta}$; iteration number from $T_{0}$ to $T_{t}$ in one epoch; batch size $B$;
1: Extract visible features $\left\{\mathbf{f}^{v}_{i}|i=1,2,\cdots,N^{v}\right\}$ and infrared features $\left\{\mathbf{f}^{r}_{i}|i=1,2,\cdots,N^{r}\right\}$;
2: Clustering features and initialize two intra-modality memory banks $\tilde{\mathbf{M}}^{v}$ and $\tilde{\mathbf{M}}^{r}$ based on the cluster centroids;
3: Do V2R MULT / R2V MULT according to Alg.1;
if $E\%2==0:$ then
       1: Initialize cross-modality memory $\mathbf{M}^{c}$ and auxiliary memory $\hat{\mathbf{M}}^{r}$ according to the results from V2R MULT;
       2: for $t=T_{0}$ to $T_{t}$ do
             Do V-based training;
else if $E\%2==1$ then
       1: Initialize cross-modality memory $\mathbf{M}^{c}$ and auxiliary memory $\hat{\mathbf{M}}^{v}$ according to the results from R2V MULT;
       2: for $t=T_{0}$ to $T_{t}$ do
             Do R-based training;
4: Output: updated network $f_{\theta}.$
Algorithm 2 Training process in one epoch.
@@
Table 1: Comparison with the state-of-the-art methods on SYSU-MM01 and RegDB. “GUR” denotes GUR without camera labels. Since our method does not require any camera label information, for fair comparison we do not report the results of GUR with camera labels.